OpenStudy (anonymous):

Solve the equation by completing the square: x2 - 6x = -4.

OpenStudy (anonymous):

now we complete the square

OpenStudy (anonymous):

half of -6 is -3 and (-3)^2=9 so write \[(x-3)^2=-4+9=5\]

OpenStudy (anonymous):

then \[x-3=\pm\sqrt{5}\] \[x=3\pm\sqrt{5}\]

OpenStudy (anonymous):

sweeet!

OpenStudy (anonymous):

and short!